OpenStudy (anonymous):

I GIVE MEDALS Create an absolute value equation that has two solutions. Explain why this equation has two solutions.

OpenStudy (anonymous):

/x-2/ = 3 solution: x - 2 = 3 and x - 2 = -3 so x = 5 and x = -1

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome.